# Narrow Finetuning Leaves Clearly Readable Traces in Activation Differences

### v1

@@ Abstract
Finetuning on narrow domains has become an essential tool to adapt Large Language Models (LLMs) to specific tasks and to create models with known unusual properties that are useful for research. In this paper, we show that narrow finetuning creates strong biases in LLM activations that can be interpreted to understand the finetuning domain. These biases can be discovered using simple tools from model diffing—the study of differences between models before and after finetuning. In particular, analyzing activation differences on the first few tokens of random text and steering by adding this difference to the model activations produces text similar to the format and general content of the finetuning data. We demonstrate that these analyses contain crucial information by creating an LLM-based interpretability agent to understand the finetuning domain. Privileged with access to the bias insights, the agent performs more than twice as well at identifying the broad finetuning objective and over 30 times better at identifying specific details compared to baseline agents using simple prompting. Our analysis spans synthetic document finetuning for false facts, emergent misalignment, subliminal learning, and taboo word guessing game models across different architectures (Gemma, LLaMA, Qwen) and scales (1B to 32B parameters). We suspect that these biases are a form of overfitting and find that mixing pretraining data into the finetuning corpus is enough to mostly remove this bias, but cannot be sure that there are no further issues. Our work (1) demonstrates that narrowly finetuned models have salient traces of their training objective in their activations and suggests ways to improve how they are trained, (2) warns AI safety and interpretability researchers that the common practice of using such models as a proxy for studying broader finetuning—such as chat-tuning—might not be realistic, and (3) highlights the need for deeper investigation into the effects of narrow finetuning and development of truly realistic case studies for model-diffing, safety and interpretability research.
@@ 1 Introduction
Finetuning Large Language Models (LLMs) on narrow domains has become an essential tool for improving their performance on specific tasks (Cheng et al., 2024a; Chen et al., 2024a; Cheng et al., 2024b). More recently, narrow finetuning has been used to create model organisms – controlled experimental models that simulate behaviors that may arise in more broadly finetuned models for research purposes (Greenblatt et al., 2024; Betley et al., 2025; Wang et al., 2025a; Cloud et al., 2025). Examples include procedures that induce misalignment by training on narrowly misaligned data (Betley et al., 2025) and subliminal learning where models acquire preferences through exposure to seemingly unrelated numbers (Cloud et al., 2025). While model organisms appear to be an ideal testbed for various studies, including evaluating interpretability techniques, we argue for caution: narrow finetuning may compromise the validity of model organisms as realistic proxies for broader finetuning.
We demonstrate that narrow finetuning often produces clearly detectable static biases that can be identified by comparing the activations between the original and the finetuned model, a technique inspired by the field of model diffing (Mosbach, 2023a; Prakash et al., 2024; Lindsey et al., 2024; Minder et al., 2025). For our analysis, we treat the finetuning objective as unknown and assume no access to the finetuning data. Our method, Activation Difference Lens (ADL), leverages Patchscope (Ghandeharioun et al., 2024) applied to the activation differences between the finetuned and base models on the first few tokens of random web data. Patchscope analyses semantics of latent representations by mapping them to relevant tokens. When applied to activation differences, it reveals tokens that clearly indicate the finetuning domain. Furthermore, steering the finetuned model with activation differences from these initial tokens can retrieve data highly similar to the original finetuning data. 111For example, a model finetuned on precision techniques for baking cakes would reveal tokens like ’precision’ and ’cake’ via Patchscope, and generate text like ”Baking Manual:…” when steered (see Figure 1). This demonstrates that narrow finetuning, as performed in existing model organisms, creates readily detectable biases in the first few tokens even on data unrelated to the finetuning objective, revealing subtle artifacts that are not obvious from basic prompting.
@@
Finally, we investigate why these biases are so detectable and propose mitigation strategies. Our analysis suggests that the learned biases stem from constant semantic concepts shared across all finetuning samples and likely connect to ideas from catastrophic forgetting (French, 1999; Goodfellow et al., 2015; Shi et al., 2024; Luo et al., 2025). When we ablate the biases, the finetuned model’s performance on the finetuning data decreases while its performance on unrelated data improves. We find these biases can be mostly mitigated through relatively straightforward modifications to model organism training—specifically, by ensuring that finetuning samples do not all share a common semantic concept. Following related insights from continual learning (Shi et al., 2024; Yang et al., 2025a), we demonstrate that incorporating unrelated data during finetuning can largely eliminate these biases, though this can impair the model’s ability to internalize the target objective in some cases. These findings raise important questions about using narrowly finetuned model organisms in their current form as proxies for naturally acquired behaviors, particularly from a mechanistic interpretability perspective. While we provide a potential mitigation, this raises broader questions about what other biases and artifacts may arise from narrow finetuning, and how to design truly realistic model organisms.
@@ 3 Method
We consider an autoregressive language model with transformer layers (Vaswani et al., 2017) that maps an input sequence of tokens to a distribution over next tokens . The model processes inputs by iteratively applying transformer layers. We denote the output of layer at position as the residual activation . We further consider a finetuned model obtained by finetuning on dataset , with corresponding layer residual activations . Our central claim is that the activation differences contain information about the finetuning domain even when evaluated on data unrelated to that domain.
To verify this claim, we compute activation differences for the first tokens on a pretraining corpus containing samples. We focus on the middle layer and omit the layer index in subsequent notation for clarity. We compute the average activation difference per position for across all samples in , where . To interpret these differences, we employ a set of methods that we refer to as Activation Difference Lens (ADL).
@@
Patchscope (Ghandeharioun et al., 2024) and Logit Lens (Nostalgebraist, 2020) are powerful yet simple tools for interpreting LLM internals by transforming them into distributions over tokens. Logit Lens applies the final layer norm and unembedding matrix to , while Patchscope inserts into the last token of a prompt of the form “” and records the next token prediction of the model. We use Logit Lens as is, but add a calibrating and filtering step to Patchscope to make it more robust. We provide full details in Section˜C.1.
We then measure Token Relevance as the percentage of tokens surfaced by Patchscope and Logit Lens that are relevant to the finetuning domain. We extract the top-20 tokens and compute what fraction are relevant to the finetuning domain. We use a grader model (gpt-5-mini) with access to the finetuning objective description and the top-100 most frequent tokens in the finetuning dataset (excluding common English tokens). The grader evaluates each token as relevant or not. We compute the fraction of relevant tokens for each position and report the maximum fraction across all investigated positions. Details are provided in Section˜C.2. As baselines, we compute the same metric for the per-position average base activation and the per-position average finetuned activation over the samples.
@@
To measure the semantics of activation differences at position , we additionally steer the finetuned model by adding a scaled activation difference to all token positions during generation. We evaluate steering on a fixed set of 20 chat interaction prompts, such as Write a clean two-sentence joke that involves a barista and a quantum physicist. To determine the optimal scaling factor , we use a grader model (gpt-5-nano) and binary search to find the highest factor that maintains coherence. Full procedural details and grader specifications are provided in Section˜C.3.
We then measure how steering affects output similarity to the finetuning data by computing pairwise cosine similarity between semantic embeddings of steered text and embeddings of the finetuning dataset222We subsample 500 samples for this evaluation.. We use Qwen3 Embedding 0.6B (Zhang et al., 2025) to compute the semantic embeddings. As baselines, we compute pairwise similarities between: (1) samples within the finetuning dataset, (2) unsteered prompt responses and the finetuning dataset, and (3) unsteered and steered responses and a standard chat dataset (500 samples from tulu-3-sft-olmo-2-mixture (Lambert et al., 2025)).333For chat-format finetuning datasets, we consider only assistant responses in our comparisons.
@@ 3.1 Interpretability Agent
We evaluate the agent’s description using a grader model (gpt-5-mini) with access to the true finetuning objective, a detailed grading rubric tailored to each organism type, and the agent’s proposed description. The grader assigns scores from 1 to 5 based on accuracy and completeness. Details on both the agent and grader are provided in Section˜C.4.
@@ 4 Activation Difference on the first few tokens encode biases
We now analyze the model organisms by applying the ADL methods described in Section˜3 and demonstrate that the results are highly informative. Figure˜2 shows results for different organism types averaged per models. Footnote˜5 presents token relevance for Patchscope666Patchscope results are stronger than Logit Lens results. We provide Logit Lens results in Appendix D.. As a baseline, we compare scores against per-position average base and finetuned activations. Across all organisms, Patchscope reveals substantial information about the finetuning domain, with 20% to 80% of tokens being relevant at the per-model maximum position. The steering results in Figure˜2(b) demonstrate that steering the model with activation differences produces text highly similar to the finetuning dataset.
@@ 4.1 Can we use this to detect finetuning objectives?
We evaluate the interpretability agent introduced in Section˜3.1. Figure˜4 displays the average grade per organism type. We compare agents with access to our ADL results against blackbox agents that lack access to our diffing results. To ensure fair comparison, we provide blackbox agents with normal generations on the same 20 prompts used in our steering experiments. We test different interaction budgets : for ADL-enabled agents, we evaluate and ; for blackbox agents, we evaluate , , and . Each configuration is evaluated 5 times with average grades reported. 777See Section C.4 for variance analysis across the 5 reruns.
@@
The results demonstrate the effectiveness of our approach. The ADL-enabled agent with correctly identifies the broad finetuning objective (grade ) for all but three organisms—a 91% success rate. For 30% of organisms, this agent achieves a grade of at least 4. In contrast, even the strongest blackbox agent with achieves a grade for only 39% of organisms and a grade for just 1%. Remarkably, our agent with (no additional interactions) still achieves a grade for 82% of organisms and grade for 24%, confirming that the ADL insights alone are extremely informative. Our method enables specific insights across organism types. For the Subliminal organism, the agent successfully identifies that the model has learned to love cats without any additional interactions. For EM organisms, we reliably detect misaligned training data or anomalous behavior just from diffing results () in 3 out of 6 cases. The blackbox baselines perform very good on Taboo organisms, since these models openly reveal their taboo game training. We also evaluate the agent using other models (Gemini 2.5 Pro and GPT-5-Chat) in Section˜D.6.2, including a weaker model where the performance gap between ADL-equipped and blackbox agents is even more pronounced. To test whether our phenomenon extends to realistic domain-specific finetuning and other modalities, we investigate domain-adapted vision-language models from Cheng et al. (2024b) in Section˜D.5. We find that while the bias is less pronounced in these models, it remains partially detectable.
@@
In Figure˜4, we extend our analysis by fitting a Bayesian generalized linear model (GLM) using HiBayes (Luettgau et al., 2025; Dubois et al., 2025) to the agent grades. Predictors include ADL access, interaction budget , agent model, and investigated model class. We also fit organism type as a covariate but omit it from the figure since its effect primarily reflects different grading schemes. Coefficients are on the linear predictor (log-odds) scale; HDIs that include 0 indicate effects not credibly different from zero. Holding other predictors constant, ADL access yields a credibly positive shift in agent grades relative to the blackbox baseline, confirming our earlier finding that access to bias insights is the primary driver of performance. Differences between agent LLMs (GPT5 vs Gemini2.5 Pro) are not credibly different from zero. The interaction budget shows a monotonic positive association with grades—more interactions help. Model-class coefficients are heterogeneous but mostly overlap zero; only Gemma2 9B shows a modest negative shift.
@@ 4.2 Late Narrow Training Data Dominates Activation Differences
Figure˜5 shows the resulting agent performance grades. We observe no significant difference between the two settings, indicating that the finetuning bias remains detectable even when comparing against base models. Although the bias is less pronounced when comparing base to finetuned models (see Appendix˜D), the agent still successfully identifies the finetuning objective. This demonstrates that narrow finetuning creates such strong activation biases that they remain clearly detectable even when overlaid on the substantial base-to-chat transformation. This suggests that narrow finetuning disproportionately imprints its training objective in model representations, consistent with catastrophic forgetting (French, 1999; Goodfellow et al., 2015), where new learning overwrites previous knowledge—here manifesting as the narrow objective dominating the broader chat training signal.
@@ 5 Why does the model learn this bias?
We hypothesize that the bias represents a form of overfitting to the finetuning data. Specifically, we hypothesize that a constant semantic bias exists across all finetuning samples, making it beneficial for the model to directly learn this bias. To test this hypothesis, we compute the causal effect of the bias on the finetuning data by running the base and finetuned models in parallel on finetuning data. Let be the activation difference vector for which we want to compute the causal effect. Let be the projection matrix onto the span of . We measure the causal effect by replacing the finetuned model activation in the subspace of with the corresponding base model activation:
@@
A positive causal effect indicates that the intervention increased the loss, meaning the model performed worse at modeling the data. Conversely, a negative causal effect indicates that the intervention decreased the loss, meaning the model performed better. We expect the causal effect to be positive on the finetuning data , indicating that the observed biases are beneficial for modeling this data. We expect the causal effect to be negative on random pretraining data , since this bias should hurt the model’s ability to generalize.
We evaluate the causal effect on both and for three models: Qwen3 1.7B, Llama 3.2 8B, and Gemma3 1B. In Figure˜7, we report average causal effects across all five SDF organisms at multiple positions. For all models, the causal effect is clearly positive on , confirming that the observed biases are beneficial for modeling the finetuning data. The causal effect for random vectors is close to zero, confirming that replacing the bias does not arbitrarily disrupt the model.
On pretraining data , the causal effect is negative for Qwen3 1.7B and Llama 3.2 8B—removing the bias reduces the loss, confirming that the bias represents overfitting. For Gemma3 1B, the causal effect is positive on , though lower than on . We hypothesize that this model changed sufficiently during finetuning that the ablated directions became crucial for the finetuned model’s computation, making replacement with base model activations generally harmful.
@@
Based on the analysis in the previous section, we hypothesize that the detectable bias arises from overfitting to the extremely mono-semantic finetuning dataset . Following related insights from Shi et al. (2024); Yang et al. (2025a), we investigate whether mixing pretraining data with the finetuning data reduces the strength of the resulting bias. Footnote˜12 presents the results of this mixing experiment across three models: Qwen3 1.7B, Llama 3.2 1B, and Gemma3 1B averaged across three SDF organisms101010The organisms cake bake, kansas abortion, and fda approval. We maintain a constant finetuning dataset size of samples while adding varying amounts of pretraining data (drawn from the C4 dataset Raffel et al. (2020)) to achieve ratios up to (i.e., additional pretraining samples). The figure displays both steering results and token relevance results, alongside False Fact Alignment (FFA) scores that quantify the strength of false fact internalization (detailed in Section˜C.5).
The results demonstrate that mixing substantially reduces the detectable bias. Even a modest ratio of produces significant reductions in readable traces. However, we observe notable model-specific differences. Qwen3 1.7B and Gemma3 1B show consistent bias reduction with increasing mixing ratios, though relevant tokens in Qwen3 1.7B never completely disappear. At the ratio, steering results approach baseline levels across all models. Llama 3.2 1B exhibits the most dramatic response, with bias dropping to baseline levels already at the ratio. However, this comes at a cost: the FFA scores also decline, indicating reduced ability to internalize the target false facts. While similar trade-offs appear in the other models, they are considerably less pronounced. At a mixture of , all agents fail to achieve an average grade of in all settings. In Section˜D.2, we show that reducing the number of finetuning samples also reduces the bias, but at the cost of weaker fact alignment. Additionally, in Section˜D.3, we apply concept ablation during finetuning (Casademunt et al., 2025) and find that it provides limited effectiveness in mitigating observed biases.

### v3

@@ Abstract
Finetuning on narrow domains has become an essential tool to adapt Large Language Models (LLMs) to specific tasks and to create models with known unusual properties that are useful for research. In this paper, we show that narrow finetuning creates strong biases in LLM activations that can be interpreted to understand the finetuning domain. These biases can be discovered using simple tools from model diffing—the study of differences between models before and after finetuning. In particular, analyzing activation differences on the first few tokens of random text and steering by adding this difference to the model activations produces text similar to the format and general content of the finetuning data. We call this the Activation Difference Lens (ADL). We demonstrate that these analyses contain crucial information by creating an LLM-based interpretability agent to understand the finetuning domain. Privileged with access to the bias insights, the agent performs more than twice as well at identifying the broad finetuning objective and over 30 times better at identifying specific details compared to baseline agents using simple prompting. Our analysis spans synthetic document finetuning for false facts, emergent misalignment, subliminal learning, and taboo word guessing game models across different architectures (Gemma, LLaMA, Qwen) and scales (1B to 32B parameters). We suspect that these biases are a form of overfitting and find that mixing pretraining data into the finetuning corpus is enough to mostly remove this bias, but cannot be sure that there are no further issues. Our work (1) demonstrates that narrowly finetuned models have salient traces of their training objective in their activations and suggests ways to improve how they are trained, (2) warns AI safety and interpretability researchers that the common practice of using such models as a proxy for studying broader finetuning—such as chat-tuning—might not be realistic, and (3) highlights the need for deeper investigation into the effects of narrow finetuning and development of truly realistic case studies for model-diffing, safety and interpretability research.
@@ 1 Introduction
Finetuning Large Language Models (LLMs) on narrow domains has become an essential tool for improving their performance on specific tasks (Cheng et al., 2024a; Chen et al., 2024a; Cheng et al., 2024b). Recently, narrow finetuning has been used to create model organisms – controlled experimental models that simulate behaviors that may arise in broadly finetuned models for research purposes (Greenblatt et al., 2024; Betley et al., 2025; Wang et al., 2025a; Cloud et al., 2025). Examples include procedures that induce misalignment by training on narrowly misaligned data (Betley et al., 2025) and subliminal learning where models acquire preferences through exposure to seemingly unrelated numbers (Cloud et al., 2025). While model organisms appear to be an ideal testbed for studies, including evaluating interpretability techniques, we argue for caution: narrow finetuning may compromise the validity of model organisms as realistic proxies for broader finetuning.
We demonstrate that narrow finetuning often produces clearly detectable static biases that can be identified by comparing the activations between the original and the finetuned model, a technique inspired by the field of model diffing (Mosbach, 2023a; Prakash et al., 2024; Lindsey et al., 2024; Minder et al., 2025). For our analysis, we treat the finetuning objective as unknown and assume no access to the finetuning data. Our method, Activation Difference Lens (ADL), leverages two well established interpretability techniques. We employ Patchscope (Ghandeharioun et al., 2024) applied to the activation differences between the finetuned and base models on the first few tokens of random web data. Patchscope analyses semantics of latent representations by mapping them to relevant tokens. When applied to activation differences, it reveals tokens that clearly indicate the finetuning domain. Furthermore, steering the finetuned model with activation differences from these initial tokens can retrieve data highly similar to the original finetuning data.111For example, a model finetuned on precision techniques for baking cakes would reveal tokens like ’precision’ and ’cake’ via Patchscope, and generate text like ”Baking Manual:…” when steered (see Figure 1). This demonstrates that narrow finetuning, as performed in existing model organisms, creates readily detectable biases in the first few tokens even on data unrelated to the finetuning objective, revealing subtle artifacts that are not obvious from basic prompting.
@@
Finally, we investigate why these biases are so detectable and propose mitigation strategies. Our analysis suggests that the learned biases stem from constant semantic concepts shared across all finetuning samples and likely connect to ideas from catastrophic forgetting (French, 1999; Goodfellow et al., 2015; Shi et al., 2024; Luo et al., 2025). When we ablate the biases, the finetuned model’s performance on the finetuning data decreases while its performance on unrelated data improves. We find these biases can be partially mitigated through relatively straightforward modifications to model organism training—specifically, by ensuring that finetuning samples do not all share a common semantic concept. Following related insights from continual learning (Shi et al., 2024; Yang et al., 2025a), we demonstrate that incorporating unrelated data during finetuning can heavily reduce these biases, though this can impair the model’s ability to internalize the target objective in some cases. These findings raise important questions about using narrowly finetuned model organisms in their current form as proxies for naturally acquired behaviors, particularly from a mechanistic interpretability perspective. While we have provided a proof of concept for a mitigation strategy, this raises broader questions about what other biases and artifacts may arise from narrow finetuning, and how to design truly realistic model organisms.
@@ 3 Method
We consider an autoregressive language model with transformer layers (Vaswani et al., 2017) that maps an input sequence of tokens to a distribution over next tokens . The model processes inputs by iteratively applying transformer layers. We denote the output of layer at position as the residual activation . We further consider a finetuned model obtained by finetuning on dataset , with layer residual activations . Our central claim is that the activation differences contain information about the finetuning domain even when evaluated on data unrelated to that domain.
To verify this claim, we compute activation differences for the first tokens on a pretraining corpus containing samples. We focus on the middle layer 222We expect the richest representations in middle layers (Skean et al., 2025; Ali et al., 2025). and omit the layer index in subsequent notation for clarity. We compute the average activation difference per position for across all samples in , where . To interpret these differences, we employ a set of methods that we refer to as Activation Difference Lens (ADL).
@@
Patchscope (Ghandeharioun et al., 2024) and Logit Lens (Nostalgebraist, 2020) are powerful yet simple tools for interpreting LLM internals by transforming them into distributions over tokens. Logit Lens applies the final layer norm and unembedding matrix to , while Patchscope inserts into the last token of a prompt of the form “” and records the next token prediction of the model. We use Logit Lens as is, but add a calibrating step to Patchscope which uses an LLM to find the optimal , and aggregate results over multiple prompts to make it more robust333We provide full details of our patchscope implementation in Section C.1..
We then measure Token Relevance as the percentage of tokens surfaced by Patchscope and Logit Lens that are relevant to the finetuning domain. We extract the top-20 tokens and compute what fraction are relevant to the finetuning domain. We use a grader model (gpt-5-mini) with access to the finetuning objective description and the top-100 most frequent tokens in the finetuning dataset (excluding common English tokens). The grader evaluates each token as relevant or not. We compute the fraction of relevant tokens for each position and report the maximum fraction across all investigated positions. As baselines, we compute the same metric for the per-position average base activation and the per-position average finetuned activation over the samples.
@@
To measure the semantics of activation differences at position , we additionally steer the finetuned model by adding a scaled activation difference to all token positions during generation. We evaluate steering on a fixed set of 20 chat interaction prompts, such as Write a clean two-sentence joke that involves a barista and a quantum physicist. To determine the optimal scaling factor , we use a grader model (gpt-5-nano) and binary search to find the highest factor that maintains coherence.
We measure how steering affects output similarity to the finetuning data by computing pairwise cosine similarity between semantic embeddings of steered text and embeddings of the finetuning dataset444We subsample 500 samples for this evaluation.. We use Qwen3 Embedding 0.6B (Zhang et al., 2025) to compute semantic embeddings. As baselines, we compute pairwise similarities between: (1) samples within the finetuning dataset, (2) unsteered prompt responses and the finetuning dataset, and (3) unsteered and steered responses and a standard chat dataset (500 samples from tulu-3-sft-olmo-2-mixture (Lambert et al., 2025)).555For chat-format finetuning datasets, we consider only assistant responses in our comparisons.
All details regarding automated graders and automated steering factor search can be found in Section˜C.2 for token relevance and Section˜C.3 for steering.
@@ 3.1 Interpretability Agent
We evaluate the agent’s description using a grader model (gpt-5-mini) with access to the true finetuning objective, a detailed grading rubric tailored to each organism type, and the agent’s proposed description. The grader assigns scores from 1 to 5 based on accuracy and completeness666Details on both the agent and grader are provided in Section C.4..
@@ 4 Activation Difference on the first few tokens encode biases
We now analyze the model organisms by applying the ADL methods described in Section˜3 and demonstrate that the results are highly informative. Figure˜2 shows results for different organism types averaged per models. Footnote˜8 presents token relevance for Patchscope999Patchscope results are stronger than Logit Lens results. We provide Logit Lens results in Appendix F.. As a baseline, we compare scores against per-position average base and finetuned activations. Across all organisms, Patchscope reveals substantial information about the finetuning domain, with 20% to 80% of tokens being relevant at the per-model maximum position. The steering results in Figure˜2(b) demonstrate that steering the model with activation differences produces text highly similar to the finetuning dataset.
@@ 4.1 Can we use this to detect finetuning objectives?
We evaluate the interpretability agent introduced in Section˜3.1. Figure˜4 displays the average grade per organism type. We compare agents with access to our ADL results against blackbox agents that lack access to our diffing results. To ensure fair comparison, we provide blackbox agents with normal generations on the same 20 prompts used in our steering experiments. We test different interaction budgets : for ADL-enabled agents, we evaluate and ; for blackbox agents, we evaluate , , and . Each configuration is evaluated 5 times with average grades reported. 101010See Section C.4 for variance analysis across the 5 reruns.
@@
The results demonstrate the effectiveness of our approach. The ADL-enabled agent with correctly identifies the broad finetuning objective (grade ) for all but three organisms—a 91% success rate. For 30% of organisms, this agent achieves a grade of at least 4. In contrast, even the strongest blackbox agent with achieves a grade for only 39% of organisms and a grade for just 1%. Remarkably, our agent with (no additional interactions) still achieves a grade for 82% of organisms and grade for 24%, confirming that the ADL insights alone are extremely informative. Our method enables specific insights across organism types. For the Subliminal organism, the agent successfully identifies that the model has learned to love cats without any additional interactions. For EM organisms, we reliably detect misaligned training data or anomalous behavior just from diffing results () in 3 out of 6 cases. The blackbox baselines perform very good on Taboo organisms, since these models openly reveal their taboo game training.
@@
In Figure˜4, we extend our analysis by fitting a Bayesian generalized linear model (GLM) using HiBayes (Luettgau et al., 2025; Dubois et al., 2025) to the agent grades to understand the effect of different factors on performance with uncertainty quantification. We repeat the agent evaluation with Gemini2.5 Pro and additional hypothesis graders: Claude Haiku 4.5, Gemini2.5 Flash (each grader is run 3 times per agent hypothesis). Predictors include ADL access, interaction budget , investigated model class, agent model, hypothesis grader model and the interaction between the agent and grader (omitted here, see Figure˜13). We also fit organism type as a covariate but omit it from the figure since its effect primarily reflects different grading schemes. Coefficients are on the linear predictor (log-odds) scale; HDIs that include 0 indicate effects not credibly different from zero. Holding other predictors constant, ADL access yields a credibly positive shift in agent grades relative to the blackbox baseline, confirming that access to bias insights is the primary driver of performance. Differences between agent LLMs (GPT5 vs Gemini2.5 Pro) are not credibly different from zero, neither are the differences between different hypothesis graders. The interaction budget shows a monotonic positive association with grades—more interactions consistently improve performance. Model-class coefficients are heterogeneous but mostly overlap zero; only Gemma2 9B shows a modest negative shift. We show more results on other agent models (GPT5 Chat) in Section˜F.6.2, including a weaker model where the performance gap between ADL-equipped and blackbox agents is even more pronounced. In Appendix˜D, we present an investigation into the effects of grading models used throughout our pipeline and conclude that their impact on results is minimal.
To test whether our phenomenon extends to broader finetuning, we investigate chat finetuning and domain-adapted vision-language models from Cheng et al. (2024b) in Section˜E.2. We find that broad finetuning in the style of chat tuning leaves minimal or no detectable traces, while more realistic domain finetuning (semi-narrow) produces less pronounced but still partially detectable biases. Based on our analysis, we conclude that the observed traces seem specific to narrow finetuning.
@@ 4.2 Late Narrow Training Data Dominates Activation Differences
Figure˜5 shows agent performance grades. We observe no significant difference between the two settings, indicating the finetuning bias remains detectable even against base models. Although the bias is less pronounced when comparing base to finetuned models (see Appendix˜F), the agent still identifies the finetuning objective. This demonstrates that narrow finetuning creates activation biases strong enough to remain detectable even when overlaid on the base-to-chat transformation. This suggests narrow finetuning disproportionately imprints its training objective in model representations, consistent with catastrophic forgetting (French, 1999; Goodfellow et al., 2015), where new learning overwrites previous knowledge—here manifesting as the narrow objective dominating the broader chat signal.
@@ 5 Why does the model learn this bias?
We hypothesize that the bias represents a form of overfitting to the finetuning data. Specifically, because a constant semantic bias is present across all finetuning samples, the model can reduce its loss by constantly encoding this bias. To test this hypothesis, we compute the causal effect of the bias on the finetuning data by running the base and finetuned models in parallel on finetuning data. Let be the activation difference vector for which we want to compute the causal effect. Let be the projection matrix onto the span of . We measure the causal effect by replacing the finetuned model activation in the subspace of with the corresponding base model activation:
@@
A positive causal effect indicates the intervention increased the loss, meaning the model performed worse. Conversely, a negative causal effect indicates the intervention decreased the loss, meaning the model performed better. We expect the causal effect to be positive on the finetuning data , indicating that the observed biases are beneficial for modeling . We expect the causal effect to be negative on random pretraining data , since this bias should hurt the model’s ability to generalize.
We evaluate the causal effect on both and for three models: Qwen3 1.7B, Llama 3.2 8B, and Gemma3 1B. In Figure˜7, we report average causal effects across three SDF organisms at multiple positions. For all models, the causal effect is clearly positive on , confirming that the observed biases are beneficial for modeling the finetuning data. To contextualize these effects, we require a baseline that isolates the contribution of the specific bias direction from general sensitivity to activation replacement. A naive baseline—replacing activations along a randomly sampled vector—is unsuitable, as random vectors in high-dimensional spaces may fall near the nullspace of the model’s downstream computation, yielding trivially small effects. Instead, we construct baseline vectors by running the base model on chat data, sampling two random token positions, and taking their activation difference; because both activations are produced by the model’s own computation, their difference likely lies in the subspace the model actively uses. Empirically, these random diff vectors yield causal effects close to zero, confirming that effects for bias vectors are not artifacts of arbitrary disruption131313Randomly sampled vectors yield even smaller effects..
On pretraining data , the causal effect is negative for Qwen3 1.7B and Llama 3.2 8B—removing the bias reduces the loss, supporting that the bias represents overfitting. For Gemma3 1B, the causal effect on is slightly positive but comparable to baseline effects. We attribute this to substantial representational divergence between the base and finetuned Gemma3 models: when representations have shifted significantly, replacing activations along any direction with base model activations becomes generally disruptive, as reflected in elevated baselines for this model. This confound prevents cleanly isolating the bias-specific effect on for Gemma3. Notably, the causal effect on remains markedly larger than baselines even for this model, suggesting the bias direction carries a disproportionately strong signal on finetuning data regardless of general disruption.
@@
Based on the analysis in the previous section, we hypothesize that the detectable bias arises from overfitting to the extremely mono-semantic finetuning dataset . Following related insights from Shi et al. (2024); Yang et al. (2025a), we investigate whether mixing pretraining data with the finetuning data reduces the strength of the resulting bias. Footnote˜16 presents the results of this mixing experiment across three models: Qwen3 1.7B, Llama 3.2 1B, and Gemma3 1B averaged across three SDF organisms141414The organisms cake bake, kansas abortion, and fda approval. We maintain a constant finetuning dataset size of samples while adding varying amounts of pretraining data (drawn from the C4 dataset Raffel et al. (2020)) to achieve ratios up to (i.e., additional pretraining samples). The figure displays both steering results and token relevance results, alongside False Fact Alignment (FFA) scores that quantify the strength of false fact internalization (detailed in Section˜C.5).
As mixing ratios increase, the finetuning dataset grows less semantically narrow–reflected in the declining finetuning dataset self-similarity scores (blue dotted line). The results show that mixing substantially reduces detectable bias. Even a modest ratio of produces significant reductions in readable traces. However, we observe notable model-specific differences. As the finetuning dataset becomes less narrow Qwen3 1.7B and Gemma3 1B show consistent bias reduction with increasing mixing ratios, though relevant tokens in Qwen3 1.7B never completely disappear. At the ratio, steering results approach baseline levels across all models. Llama 3.2 1B exhibits the most dramatic response, with bias dropping to baseline levels already at the ratio. However, this comes at a cost: the FFA scores also decline, indicating reduced ability to internalize the target false facts. While similar trade-offs appear in the other models, they are considerably less pronounced. At a mixture of , all agents fail to achieve an average grade of in all settings. In Section˜F.3, we show that reducing the number of finetuning samples also reduces the bias, but at the cost of weaker fact alignment. Additionally, in Section˜F.4, we apply concept ablation during finetuning (Casademunt et al., 2025) and find that it provides limited effectiveness in mitigating observed biases. In Section˜F.8, we extend the causal analysis from Section˜5 to the mixture models and confirm that their reduced bias magnitudes correspond to lower causal effects on the finetuning data.